NACE EXAM — Health Care exam material
HEAL100 · Other
HEAL100 Health Care — exam and revision material · Other
Updated today
Health Care exam material, 31 pages. Originally published on StudyMerit.
See the first 3 pages exactly as they appear in the document.
Stuvia.com - The Marketplace to Buy and Sell your Study Material Downloaded by: patriciac | chevylove4ford2@gmail.com Distribution of this document is illegal Want to earn $1.236 extra per year? C784 / Objective Assessment (Latest Update 2025 / 2026) Applied Healthcare Statistics | Module 7| Questions and Answers | Grade A | 100% Correct – WG 1. What is the sample space for choosing an ice cream flavor (chocolate, strawberry, vanilla) and sauce (chocolate, caramel, or none)? ANS The Answer is 9 . From the list that was constructed, there are 9 outcomes. Vanilla, Chocolate Vanilla, Caramel Vanilla, None Chocolate, Chocolate Chocolate, Caramel Chocolate, None Strawberry, Chocolate Strawberry, Caramel Strawberry, None 1
Stuvia.com - The Marketplace to Buy and Sell your Study Material Downloaded by: patriciac | chevylove4ford2@gmail.com Distribution of this document is illegal Want to earn $1.236 extra per year? 2. In another game, you win if you roll a number less than 3 or if you roll an even number. What is your probability of winning this game? ANS P(A or B) = P(A) + P(B) - P(A and B). There are 2 numbers less than 3, so the probability of event A is 2/6 or 1/3. There are 3 numbers that are even, so P(B) = 1/2. But the number 2 is both less than 3 AND even, so P(A and B) = 1/6. The formula gives us 1/3 + 1/2 - 1/6 = 2/3. 3. What is the sample space for flipping 3 coins? ANS The Answer is 8 . From the list constructed below, there are 8 outcomes. Heads, Heads, Heads Heads, Heads, Tails Heads, Tails, Heads Heads, Tails, Tails Tails, Tails, Tails Tails, Tails, Heads Tails, Heads, Tails Tails, Heads, Heads 4. A frequency table consists of two categorical variables. True or False? ANS TRUE A frequency (or two-way) table contains two different categorical variables. 2
Stuvia.com - The Marketplace to Buy and Sell your Study Material Downloaded by: patriciac | chevylove4ford2@gmail.com Distribution of this document is illegal Want to earn $1.236 extra per year? 5. If the probability of a sunny day is 0.7 and the probability of the date on the calendar being an odd number is 0.5, what is the probability of the calendar date being odd or it being a sunny day? ANS P(A or B) = P(A) + P(B) - P(A and B ). Since "calendar date being an odd number" and "being a sunny day" are independent events, P(A and B) = 0.7 * 0.5 = 0.35. Calendar date being an odd numbered day or being a sunny day = 0.7 + 0.5 - 0.35 = 0.85. 85% chance of it being sunny or being an odd numbered day on the calendar. 6. The sum of all probabilities of individual outcomes in the sample space must equal 1. True or False? ANS TRUE The sum of all probabilities of individual outcomes equals 1. 7. You roll a six-sided die twice. What is the probability of rolling two sixes? ANS 1/36 The size of the sample space of the possible outcomes after two rolls is 36. These outcomes are equally likely. There is one outcome that results in two sixes. 8. A number shown on a die is even and a number is less than 3. disjoint or non- disjoint? ANS NON DISJOINT 2 is even and 2 is less than 3, so these events are not disjoint. 3
28 more pages after this
You’ve read 3 of 31. Unlock once and it stays in your library — and if the rest doesn’t match what you just read, you get a refund.
HEAL100 · Other
HEAL100 · Other
HEAL100 · Other
$10.00
Yours permanently